OpenStudy (anonymous):

A motorcycle cost $12,000 when it was purchased. The value of a motorcycle decreases by 6% each year. Find the rate of decay each month and select the correct answer below.

OpenStudy (anonymous):

@Michele_Laino @mathmate @phi

OpenStudy (anonymous):

1 year has 12 months so whats 6 divide by 12?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

each month it depreciates by 0.5%

OpenStudy (anonymous):

I need one more:)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You decide to put $175 in a savings account to save for a $3,000 down payment on a new car. If the account has an interest rate of 3% per year and is compounded monthly, how long does it take you to earn $3,000 without depositing any additional funds?

OpenStudy (anonymous):

Compound interest formula\[A=P(1+\frac{ r }{ 100 })^n\]\[3000=175(1+0.3)^n\]

OpenStudy (anonymous):

300=227.5^n

OpenStudy (anonymous):

no it doesnt work like this

OpenStudy (anonymous):

\[17.14285714=1.03^n\]

OpenStudy (anonymous):

16.64^n

OpenStudy (anonymous):

?

OpenStudy (anonymous):

log 17.14285714 = 1.23408320596 log 5 = 0.69897000433 1.23408320596/0.069897000433

OpenStudy (anonymous):

Thats not an answer choice 8.0111 years 96.1332 years 93.2914 years 94.8377 years

OpenStudy (anonymous):

plug in all the in answers to the compound interest formula and try out

OpenStudy (anonymous):

? What do you mean?

OpenStudy (anonymous):

175(1.03)^n replace n with all the 4 options, see which one works

OpenStudy (mathmate):

or you take log!

OpenStudy (anonymous):

i did

OpenStudy (phi):

interest rate of 3% per year and is compounded monthly, that means you work in months (not years) and the interest rate is 0.03/12 the equation you want to solve is 175(1+ 0.03/12)^n = 3000 175(1.0025)^n = 3000 1.0025^n= 17.142857

OpenStudy (anonymous):

dayum i forgot something...the interest is compounded monthly!

OpenStudy (mathmate):

1.0025^n=3000/175 n log(1.0025)=log(3000/175) solve for n.

OpenStudy (anonymous):

knew something was off

OpenStudy (anonymous):

Ok hold on:)

OpenStudy (anonymous):

For some reason I got C?? I think I did something wrong

OpenStudy (phi):

when you find n, it will be in months divide by 12 to get years

OpenStudy (anonymous):

Was I right?

OpenStudy (phi):

what do you get for log(3000/175) ?

OpenStudy (anonymous):

1.23408320603

OpenStudy (phi):

so you have n log(1.0025) = 1.23408320603 divide both sides by log(1.0025) n= 1.23408320603/log(1.0025) what do you get for n

OpenStudy (anonymous):

1.2310056918=n

OpenStudy (anonymous):

1138.05283703 i mean

OpenStudy (phi):

ok, and that is in months divide by 12 to get years

OpenStudy (anonymous):

94.8377...

OpenStudy (phi):

yes

OpenStudy (anonymous):

ok so going back to #1 i think sesal gave me the wrong answer??

OpenStudy (anonymous):

A motorcycle cost $12,000 when it was purchased. The value of a motorcycle decreases by 6% each year. Find the rate of decay each month and select the correct answer below.

OpenStudy (anonymous):

I got -0.005143% as an swer and she got -0.5% so I dont even know

OpenStudy (phi):

if it has a rate of 6% per year, or in this case -6%, as a decimal is -0.06 on a monthly basis, divide by 12: -0.06/12 = -0.005 in percent: -0.5%

OpenStudy (anonymous):

Ok! Thanks :) I was just checking!

OpenStudy (anonymous):

Solve 3^(2x)=7^(x-1) I got -7.74293 ??

OpenStudy (phi):

if we put your answer into the equation it works

OpenStudy (anonymous):

Thanks:)